# From Release to Adoption: Challenges in Reusing Pre-trained AI Models for Downstream Developers

### v1

@@ I Introduction
However, reusing PTMs within software systems introduces new forms of complexity beyond traditional software dependencies [7]. PTMs are often treated as opaque components, with minimal documentation [3], unclear behavior [8], lack of standardized versioning [9], and significant computational demands [10]. As a result, developers who are not involved in model training or architecture design often encounter difficulties in using models for their purposes [11, 12, 13]. We refer to these developers as downstream developers—that is, practitioners who reuse PTMs in their own software (i.e., downstream software) without directly contributing to the model’s development. These developers differ from model creators or ML researchers; they primarily work in software engineering roles and must adapt PTMs to fit within their downstream software environments and constraints.
@@ III-B Data Collection and Pre-processing
Filtering Software Projects: To ensure the quality and relevance of data, we filtered the dataset in the following steps: Firstly, aiming to investigate the challenges of commonly-used PTMs, we included PTMs with more than 2K downloads as suggested by previous studies [7, 34, 21, 35], reducing the dataset to 5,938 projects and 197,882 issues. Secondly, we applied GitHub-based metrics, as seen in Figure 2, to ensure the selection of active, well-maintained, and relevant projects. These include size, popularity, activity, data availability, content filtering (excluding tutorials, homework, coding challenges, or resource collections via keyword-based filtering from [36]), issues, and commits. These criteria are adapted from best practices in software repository analysis [37, 38, 36, 11, 39, 40]. Notably, the thresholds in both steps were chosen based on elbow points, where the curve levels off when plotting each metric against the number of remaining projects. This resulted in 293 projects and 116,658 issues. Then, we included only projects written in Python, as it is the most popular programming language for AI-related projects [37, 41] and the predominant language in the dataset, yielding 253 projects and 101,596 issues. After that, we manually inspected and excluded projects where most content was in a non-English language, such as Bark-Voice-Cloning [^2] https://github.com/KevinWang676/Bark-Voice-Cloning, resulting in 227 projects and 93,061 issues.
@@ III-C Data Analysis
To further mitigate potential confounding effects, particularly those stemming from differences in issue types (e.g., bug, question, enhancement), we used a fine-tuned seBERT model [46] to classify the issues into these categories. It is important to note that this classification was used solely to provide a high-level overview of the issue type distribution and to validate the balance between the groups. To strengthen the validity of this classification, we conducted a broader manual validation. Using a statistically representative sample of 260 issues (130 PTM and 130 non-PTM) with 95% confidence level and 5% margin of error, we found that 249 cases (95.77%) matched the manual classification, confirming the reliability of the seBERT-based issue type tagging. A Chi-square independence test [47] confirmed that the distribution of issue types did not significantly differ between the groups (p-value = 0.2), supporting the comparability of the two sets.
@@ IV-A Model Usage
IV-B Model Performance and Output Quality
@@ IV-C Software Environment
IV-D Developer Information and Documentation
@@
IV-E Hardware and Computation Resource
@@
IV-F Model Resource and Configuration
@@
IV-G Model Functionality Enhancement Needs
@@ V Resolution Time (RQ2)
Figure 5(b) presents the average and median resolution times across top-level challenge themes. A Kruskal-Wallis H-test reveals a statistically significant difference in resolution time across different challenge categories ($p<0.001$). This indicates that certain challenge categories take disproportionately longer to resolve than others. Model Performance and Output Quality has the highest average resolution time (79 days). The longest sub-theme is Model Training and Evaluation, averaging 202 days. Within this sub-theme, Model Training Needs, which includes requirements for model retraining and fine-tuning to improve accuracy, was the most prolonged issue. Note that only the top-level themes are shown in Figure 5(b).
The findings also show that Model Functionality Enhancement Needs has the second-highest average resolution time, at 48 days. We found that issues under this theme often involve enhancements rather than traditional bug fixes. These issues require developers to assess multiple factors before deciding to implement them, such as implementation complexity, availability of existing models or functionalities, project scope limitations, accuracy concerns, potential risks of misuse, and development costs. Notably, we observed multiple cases where feature requests were rejected outright, not due to technical infeasibility, but rather because they failed to meet the above considerations, as exemplified in babyagi #105[98], imaginAIry #116[99], and imaginAIry #132[100].
@@ VIII-C Construct Validity
Our study aims to identify challenges specifically related to the reuse of PTMs in downstream software projects. However, there is a risk that a few identified sub-themes may appear generic to general software systems (e.g., GUI). To mitigate this, we focus our analysis on how these challenges manifest themselves uniquely in the context of PTM reuse. In RQ2, resolution time may not always reflect actual fix effort, as compounding factors such as issue type, developer workload, and limited interaction with the reporter may influence it [40]. To mitigate this threat, we controlled the distribution for two significant factors—project and issue type—by employing stratified random sampling when selecting non-PTM issues to match the PTM group. Both groups were drawn from the same annotated dataset, and the matching process ensured a balanced distribution across projects and issue types. To validate the reliability of the issue type classification by seBERT used for stratification, we also performed a broader manual validation of a representative sample to indicate that the use of seBERT introduces minimal risk of systematic bias. While this reduces sampling bias and improves comparability, we acknowledge that residual confounding factors and sensitivity to sample variation may still influence results. We also accounted for extreme outliers by using non-parametric tests, ensuring our analysis is less sensitive to anomalies.

### v2

@@ I Introduction
However, reusing PTMs within software systems introduces new forms of complexity beyond traditional software dependencies [7]. PTMs are often treated as opaque components, with minimal documentation [3], unclear behavior [8], lack of standardized versioning [9], and significant computational demands [10]. As a result, developers who are not involved in model training or architecture design often encounter difficulties in using models for their purposes [11, 12, 13]. We refer to these developers as downstream developers—that is, practitioners who reuse PTMs in their own software (i.e., downstream software) without directly contributing to the model’s development. These developers differ from model creators or ML researchers; they primarily work in software engineering roles and adapt PTMs to fit within their downstream software environments and constraints.
@@ III-B Data Collection and Pre-processing
Filtering Software Projects: To ensure the quality and relevance of data, we filtered the dataset in the following steps: Firstly, aiming to investigate the challenges of commonly-used PTMs, we included PTMs with more than 2,000 downloads as suggested by previous studies [7, 34, 21, 35], reducing the dataset to 5,938 projects and 197,882 issues. Secondly, we applied GitHub-based metrics, as seen in Figure 2, to ensure the selection of active, well-maintained, and relevant projects. These include size, popularity, activity, data availability, content filtering (excluding tutorials, homework, coding challenges, or resource collections via keyword-based filtering from [36]), issues, and commits. These criteria are adapted from best practices in software repository analysis [37, 38, 36, 11, 39, 40]. Notably, the thresholds in both steps were chosen based on elbow points, where the curve levels off when plotting each metric against the number of remaining projects. This resulted in 293 projects and 116,658 issues. Then, we included only projects written in Python, as it is the most popular programming language for AI-related projects [37, 41] and the predominant language in the dataset, yielding 253 projects and 101,596 issues. After that, we manually inspected and excluded projects where most content was in a non-English language, such as Bark-Voice-Cloning [^2] https://github.com/KevinWang676/Bark-Voice-Cloning, resulting in 227 projects and 93,061 issues.
@@ III-C Data Analysis
To further mitigate potential confounding effects, particularly those stemming from differences in issue types, we used a fine-tuned seBERT model [46] to classify issues into three categories: bug, question, and enhancement. These are the predefined issue types supported by the model, and the classification was used solely to provide a high-level overview of issue type distribution and validate the balance between the PTM and non-PTM groups. To strengthen the validity of this classification, we conducted a broader manual validation. Using a statistically representative sample of 260 issues (130 PTM and 130 non-PTM) with 95% confidence level and 5% margin of error, we found that 249 cases (95.77%) matched the manual classification, confirming the reliability of the seBERT-based issue type tagging. A Chi-square independence test [47] confirmed that the distribution of issue types did not significantly differ between the groups (p-value = 0.2), supporting the comparability of the two sets.
@@ IV-A Model Usage (MU)
IV-B Model Performance and Output Quality (PQ)
@@ IV-C Software Environment (SE)
IV-D Developer Information and Documentation (DI)
@@
IV-E Hardware and Computation Resource (HR)
@@
IV-F Model Resource and Configuration (RC)
@@
IV-G Model Functionality Enhancement Needs (FN)
@@ V Resolution Time (RQ2)
Figure 5(b) presents the average and median resolution times across top-level challenge themes. A Kruskal-Wallis H-test reveals a statistically significant difference in resolution time across different challenge categories ($p<0.001$). This indicates that certain challenge categories take disproportionately longer to resolve than others. Model Performance and Output Quality (PQ) has the highest average resolution time (79 days). The longest sub-theme is Model Training and Evaluation, averaging 202 days. Within this sub-theme, Model Training Needs, which includes requirements for model retraining and fine-tuning to improve accuracy, was the most prolonged issue. Note that only the top-level themes are shown in Figure 5(b).
The findings also show that Model Functionality Enhancement Needs (FN) has the second-highest average resolution time, at 48 days. We found that issues under this theme often involve enhancements rather than traditional bug fixes. These issues require developers to assess multiple factors before deciding to implement them, such as implementation complexity, availability of existing models or functionalities, project scope limitations, accuracy concerns, potential risks of misuse, and development costs. Notably, we observed multiple cases where feature requests were rejected outright, not due to technical infeasibility, but rather because they failed to meet the above considerations, as exemplified in babyagi #105[98], imaginAIry #116[99], and imaginAIry #132[100].
@@ VIII-C Construct Validity
Our study aims to identify challenges specifically related to the reuse of PTMs in downstream software projects. However, there is a risk that a few identified sub-themes may appear generic to general software systems (e.g., GUI). To mitigate this threat, we focus our analysis on how these challenges manifest themselves uniquely in the context of PTM reuse. In RQ2, resolution time may not always reflect actual fix effort, as compounding factors such as issue type, developer workload, and limited interaction with the reporter may influence it [40]. To mitigate this threat, we controlled the distribution for two significant factors—project and issue type—by employing stratified random sampling when selecting non-PTM issues to match the PTM group. Both groups were drawn from the same annotated dataset, and the matching process ensured a balanced distribution across projects and issue types. To validate the reliability of the issue type classification by seBERT used for stratification, we also performed a broader manual validation of a representative sample to indicate that the use of seBERT introduces minimal risk of systematic bias. While this reduces sampling bias and improves comparability, we acknowledge that residual confounding factors and sensitivity to sample variation may still influence results. We also accounted for extreme outliers by using non-parametric tests, ensuring our analysis is less sensitive to anomalies.